# Video Frame Interpolation with Region-Distinguishable Priors from SAM

### v1

@@ 4.5 User Study
Fig. 8 displays the results of the user study, revealing that our method received more selections from participants compared to all the baselines. While some participants opted for the ”same” option, this is primarily attributed to the resolution of the testing images. Higher resolution tends to amplify differences, as observed in the results from the SNU-FILM dataset. This underscores that our method can enhance the human subjective perception of baselines.
5 Limitations
While our proposed method has achieved commendable performance improvement on multiple datasets, there are several limitations that we aim to address in future work. First, we plan to investigate more lightweight approaches, such as employing advanced networks to further reduce the parameter and computation cost. Additionally, we will explore strategies that consistently yield further improvements across all benchmarks.
6 Conclusion
In this work, we introduced a plug-and-play module designed to enhance the performance of existing VFI approaches. We innovatively designed RDPs using SAM and implemented the HRFFM to integrate them into VFI methods. Extensive experiments demonstrate that our strategy significantly improves the performance of current VFI methods, achieving SOTA results across multiple well-recognized benchmarks.

### v2

@@ IV-F User Study
Fig. 9 displays the results of the user study, revealing that our method received more selections from participants compared to all the baselines. While some participants opted for the “same” option, this is primarily attributed to the resolution of the testing images. Higher resolution tends to amplify differences, as observed in the results from the SNU-FILM dataset. This underscores that our method can enhance the human subjective perception of baselines.
V Limitations
While our proposed method has achieved commendable performance improvement on multiple datasets, there are several limitations that we aim to address in future work. First, we plan to investigate more lightweight approaches, such as employing advanced networks to further reduce the parameter and computation cost.
VI Conclusion
In this work, we introduced a plug-and-play module designed to enhance the performance of existing VFI approaches. We innovatively designed RDPs using SAM2 and implemented the HRFFM to integrate them into VFI methods. Extensive experiments demonstrate that our strategy significantly improves the performance of current VFI methods, achieving SOTA results across multiple well-recognized benchmarks.